OpenStudy (anonymous):

In a simple random sample of 219 students at a college, 73 reported that they have at least $1000 of credit card debt. What is the 99% confidence interval for the percent of all the students at that college who have at least $1000 in credit card debt? (Points : 3) (25.0, 41.6) (27.5, 39.1) (30.1, 36.5) (31.0, 35.6)

OpenStudy (anonymous):

Take note of the formula for the confidence interval for population proportions.\[\displaystyle \hat{p} \pm \text{z}_{\alpha/2} * (\sqrt{\displaystyle \frac{\hat{p} * \hat{q}}{n}})\]

OpenStudy (anonymous):

Where \[\hat p\] is the sample proportion, we apply the formula: \[\hat p = \frac{x}{n}\] where x is the number of successes and n is the sample size

OpenStudy (anonymous):

The \[z_{\alpha/2}\] can be found using the z-table.

OpenStudy (anonymous):

Let me attached an example on finding the \[z_{\alpha/2}\].

OpenStudy (anonymous):

Suppose we have a confidence level of 90%. Find the value of \[\alpha/2\] first. Apply the formula: \[\frac{\alpha}{2} = \frac{1-c}{2}\] where c is the confidence level written in decimal form. So, for this example, c = 0.90. \[\frac{\alpha}{2} = \frac{1-0.90}{2}= 0.05\]

OpenStudy (anonymous):

We will locate the 0.05 on the z-table. The 0.05 is in between 0.0495 and 0.0505. So, we will take the average of the z-scores that correspond to 0.0495 and 0.0505.

OpenStudy (anonymous):

We remove the negative sign. \[z_{a/2}= \frac{1.64 + 1.65}{2} = 1.645\]